OpenStudy (anonymous):

Given the following rational function: (a) state the domain. (b) find the vertical and horizontal asymptotes, if any. (c) find the oblique asymptotes, if any. (d) submit a graph of this function to the Dropbox. Please show all of your work. f(x)=(x^2+ 2x-8)/(x^2 + 7x + 12)

OpenStudy (lgbasallote):

Domain: all values of x except -4 and -3

OpenStudy (anonymous):

can you show me the steps on how you got that ?

Directrix (directrix):

f(x)=(x^2+ 2x-8)/(x^2 + 7x + 12) factors to [(x-2 )(x+4)] / {(x + 3) ( x + 4)] = f(x)

Directrix (directrix):

The horizontal asymptote has equation y = 1. As x increases without bound, y = (x-2)/(x+3) approaches 1.

Directrix (directrix):

There may be just one vertical asymptote which would be x = -3. [(x-2 )(x+4)] / {(x + 3) ( x + 4)] = f(x) The (x+4) terms divide out giving (x - 2) / (x + 3) = f(x). The "problem" at x = -4 may be a removeable discontinuity. See graph at: http://www.wolframalpha.com/input/?i=y+%3D%28x%5E2%2B+2x-8%29%2F%28x%5E2+%2B+7x+%2B+12%29 while I read check on removeable discontinuties.

Directrix (directrix):

Removeable Discontinuity at (-4,6) --> shows as "hole" on graph y-intercept (0,-2/3) x-intercept (2,0) vertical asymptote x = -3 horizontal asymptote y = 1 no oblique asymptote  Removeable Discontinuity at (-4,6)